OpenStudy (anonymous):

ok 3/y-1 + 5/4y I just need help to figure out the LCM of the denom

OpenStudy (anonymous):

\[5y(y-1)\]

OpenStudy (anonymous):

that's the answer or the LCM?

OpenStudy (anonymous):

lcm

OpenStudy (anonymous):

how did you figure that out? is there a trick?

OpenStudy (anonymous):

\[\frac{3}{y-1}+\frac{5}{4y}=\frac{3(4y)+5(y-1)}{y-1}\] \[=\frac{12+5y-5}{4y(y-1)}=\frac{5y+7}{4y(y-1)}\]

OpenStudy (anonymous):

no trick, if there are no common factors you have to multiply the denominators together.

OpenStudy (anonymous):

just like if i had \[\frac{3}{5}+\frac{5}{11}\] i would have to multiply the denominators together

OpenStudy (anonymous):

ok well thank you:)

OpenStudy (anonymous):

yw